# Decoding Neural Emotion Patterns through Large Language Model Embeddings

### v2

@@ 3.3 Experiment 3: Human versus LLM Chatbot
The third experiment was performed on the Schema-Guided Dialogue dataset [40], which represents nearly half a million sentences comprised of human and LLM chatbot interactions. Comparing human conversational texts with LLM-generated responses revealed systematic divergences in neural activation profiles across limbic, cortical, and brainstem regions. Figure 8 shows a 3D cortical rendering in MNI space, with lateral (top) and dorsal (bottom) views indicating the magnitude of differential activation between human subjects (left) and LLM chatbot (right) responses. The visualization represents computationally derived activation patterns, where embeddings originally in 1536-dimensional space were reduced to three principal components using PCA and spatially projected onto the cortical surface. Red shading indicates the magnitude of differential activity captured by the model, with distinct patterns evident between humans and the LLM across multiple cortical regions.
Figure 8: 3D rendering of activation differences (Table 5) showing lateral (top) and dorsal (bottom) views between human subjects (left) and an LLM chatbot (right).
@@
Humans demonstrated stronger recruitment of emotion-related regions, including bilateral amygdalae [79, 80], as well as memory-related structures such as the left hippocampus, consistent with reliance on autobiographical retrieval during dialogue [81]. The human text also showed greater engagement of reward and arousal-related circuits, including the ventral tegmental area, raphe nuclei, and locus coeruleus, reflecting dopaminergic and serotonergic modulation of motivation and adaptive arousal [82, 83]. Increased activity in medial prefrontal cortex, posterior cingulate, and temporal poles further supports integration of self-referential and affective context into human conversation [84, 85].
In contrast, LLM-generated responses showed heightened anterior cingulate activity bilaterally, aligning with its role in monitoring and conflict regulation [86, 49]. LLMs also engaged the right hippocampus more strongly, suggesting an episodic-associative rather than autobiographical memory profile [81]. Cortical valuation and decision-making appeared lateralized, with left orbitofrontal cortex stronger in LLMs, while right orbitofrontal cortex was greater in humans [87, 88]. Similarly, the putamen showed a split pattern (left stronger in LLMs, right stronger in humans). The superior temporal gyri was not stronger in LLMs. Instead, the left was more active in humans, consistent with its role in speech and semantic processing [89, 90]. Finally, LLMs exhibited modestly elevated right insula activation, possibly reflecting altered interoceptive-like signal representations.
Together, these findings indicate that humans preferentially engage limbic brainstem networks integrating affect, memory, and motivation into language use, whereas LLMs display a bias toward cingulate, orbitofrontal, and striatal pathways associated with conflict monitoring and associative sequencing. This dissociation is consistent with recent work contrasting artificial and biological language networks [10, 11].
Our proposed approach therefore shows promise in distinguishing human-authored text from LLM-generated content, supporting recent studies [32, 31, 29, 30] that have demonstrated the potential of computational approaches in analyzing text to predict and classify various characteristics. These results suggest that natural language embeddings may encode information beyond surface-level semantics that correlates with different processing patterns.
@@ 4 Study Limitations
Several important limitations must be acknowledged. Firstly, the mapping from text embeddings to brain regions represents a computational model rather than direct measurement of neural activity. Recent advances in computational neuroscience have demonstrated the feasibility of embedding-to-brain mappings: for example, Goldstein et al. [26] established that linear encoding models can effectively predict neural activity from high-dimensional text embeddings, achieving correlations up to 0.50 with actual neural signals during natural conversations. While this provides methodological support for computational approaches, our study does not directly validate these mappings against neuroimaging data.
Secondly, although brain signals from fMRI and EEG can be decoded into coherent text [91, 92, 93], the inverse problem of predicting potentially engaged brain regions during language processing based on text input remains largely unexplored and requires careful validation. Goldstein et al. [26] highlighted both the promise and the challenge of this approach, demonstrating systematic individual and regional differences in how embedding-based models align with neural activity. These findings underscore the importance of extending our population-level predictions to individual-level validation across different modalities.
Finally, the predefined regional coordinates used in our study, while grounded in established neuroimaging research, represent population averages that may not accurately capture individual neuro-anatomy or the mixed selectivity patterns revealed in prior work [26]. Future studies should address these limitations by combining embedding-based predictions with direct neuroimaging measures and by accounting for individual variability in brain-language mappings, particularly in atypical or clinical populations.
@@ 5 Conclusion
In this study, we developed a novel decoding scheme that maps neural emotion patterns to brain regions through LLM embeddings. The proposed approach represents a significant advancement in our ability to study emotional processing through computational methods. By leveraging state-of-the-art natural language processing techniques and established neuro-anatomical knowledge, the proposed approach offers a scalable, accessible alternative to traditional neuroimaging methods. Our proposed approach’s ability to process natural language and map emotional content onto anatomically defined brain regions opens new possibilities for understanding individual differences in emotional processing, monitoring mental health at scale, and developing personalized interventions.
While important limitations exist, the approach offers compelling advantages in terms of cost, accessibility, and ecological validity. Future research should focus on validating the approach against established neuroimaging methods, addressing methodological limitations and exploring novel applications in clinical and research contexts. The integration of this computational approach with traditional neuro-scientific methods has the potential to accelerate our understanding of the neural basis of emotion and contribute to more effective treatments for emotional and psychiatric disorders. To encourage further exploration and application of the proposed approach, the complete source code used in this study is publicly available on GitHub at: https://github.com/xalentis/EmotionBrainMapping.

### v3

@@ 3.3 Experiment 3: Human versus LLM Chatbot
The third experiment was performed on the Schema-Guided Dialogue dataset dataset_google , which represents nearly half a million sentences comprised of human and LLM chatbot interactions. Comparing human conversational texts with LLM-generated responses revealed systematic divergences in predicted activation profiles across limbic, cortical, and brainstem regions.
Figure 8 shows a 3D cortical rendering in MNI space, with lateral (top) and dorsal (bottom) views indicating the magnitude of differential activation between human subjects (left) and LLM chatbot (right) responses. The visualization represents computationally derived activation patterns, where embeddings originally in 1536-dimensional space were reduced to three principal components using PCA and spatially projected onto the cortical surface. Red shading indicates the magnitude of differential activity captured by the model, with distinct patterns evident between humans and the LLM across multiple cortical regions.
Figure 8: 3D rendering of computational activation maps (Table 5) showing lateral (top) and dorsal (bottom) views between human subjects (left) and an LLM chatbot (right). The color bar indicates normalized activation magnitudes, ranging from 0.00 (white) to 0.50 (red).
@@
Humans demonstrated stronger recruitment of emotion-related regions, including bilateral amygdalae phelps2006 , adolphs2013 , as well as memory-related structures such as the left hippocampus, consistent with reliance on autobiographical retrieval during dialogue moscovitch2016 . The human text also showed greater engagement of reward and arousal-related circuits, including the ventral tegmental area and raphe nuclei, reflecting dopaminergic and serotonergic modulation of motivation and adaptive arousal aston2012 , dayan2012 . Greater engagement of the posterior cingulate and temporal poles further supports the integration of self-referential and affective context into human conversation raichle2001 , amodio2006 .
In contrast, LLM-generated responses showed heightened anterior cingulate activity bilaterally, aligning with its role in monitoring and conflict regulation botvinick2004 , etkin2011 . LLMs also engaged the right hippocampus more strongly, suggesting an episodic-associative rather than autobiographical memory profile moscovitch2016 . Cortical valuation and decision-making appeared lateralized, with left orbitofrontal cortex stronger in LLMs, while right orbitofrontal cortex was greater in humans rolls2004 , schoenbaum2009 . Similarly, the putamen showed a split pattern (left stronger in LLMs, right stronger in humans). The superior temporal gyri was not stronger in LLMs. Instead, the left was more active in humans, consistent with its role in speech and semantic processing hickok2007 , price2012 . Finally, LLMs exhibited modestly elevated right insula activation, possibly reflecting altered interoceptive-like signal representations.
Together, these findings indicate that humans preferentially engage limbic brainstem networks integrating affect, memory, and motivation into language use, whereas LLMs display a bias toward cingulate, orbitofrontal, and striatal pathways associated with conflict monitoring and associative sequencing. This dissociation is consistent with recent work contrasting artificial and biological language networks caucheteux2022 , toneva2022 .
Our proposed approach therefore shows promise in distinguishing human-authored text from LLM-generated content, supporting recent studies zhong2025 , lorenzoni2025 , liu2025 , ge2025 that have demonstrated the potential of computational approaches in analyzing text to predict and classify various characteristics. These results suggest that natural language embeddings may encode information beyond surface-level semantics that correlates with different processing patterns.
An important limitation of this experiment is that we compare human-authored conversational turns with LLM-generated responses to those same human utterances, rather than comparing independent human-to-human versus LLM-to-LLM dialogues. Consequently, observed differences may reflect response versus initiation dynamics rather than fundamental differences in emotional expression capacity. Furthermore, as language models continue to evolve with improvements in contextual understanding, emotional nuance, and conversational naturalness, the specific patterns we observe may shift substantially.
@@ 4 Study Limitations
Several important limitations must be acknowledged. First, the mappings from embeddings to brain regions are computational inferences, not direct measures of neural activity. Although recent work by Goldstein et al. Goldstein2025 shows that language-derived embeddings can partially predict neural responses, our approach has not been validated against imaging data and should therefore be viewed as hypothesis-generating rather than confirmatory.
Second, while brain signals can now be decoded into coherent text from fMRI and EEG recordings qiu2025 , tang2023 , levy2025 , the inverse process of predicting likely regional activation from text remains largely exploratory. The current results rely on population-average coordinates and thus do not capture individual anatomical variability or mixed-selectivity patterns observed in prior neuroimaging research Goldstein2025 .
Finally, because the regional coordinates were predefined from meta-analytic studies, the observed correspondences partly reflect built-in anatomical constraints. Future work should integrate embedding-based predictions with empirical neuroimaging across individuals and modalities to test these computationally derived hypotheses.
@@ 5 Conclusion
This study introduces a scalable computational framework that links emotional language to anatomically defined brain regions using embedding-based representations. By combining natural language processing with established neuro-anatomical knowledge, the method provides a cost-effective and interpretable complement to traditional neuroimaging approaches.
Across three experiments, the framework differentiated between healthy and depressed language patterns, characterized emotion-specific activation hierarchies, and revealed systematic contrasts between human and LLM-generated text. These results demonstrate the feasibility of embedding-to-brain mapping as a tool for generating testable hypotheses about emotional processing.
While the approach requires empirical validation, its ability to model affective variability directly from text suggests potential applications in scalable mental-health assessment and neuro-computational research. Future studies should focus on integrating this computational method with imaging data to evaluate its predictive validity and refine its neuro-biological grounding. To encourage further exploration and application of the proposed approach, the complete source code used in this study is publicly available on GitHub at: https://github.com/xalentis/EmotionBrainMapping.